OpenStudy (anonymous):

91√2t+5t1√28t−3t4√8t

OpenStudy (turingtest):

USE PARENTHESES!!!!!!!!!

OpenStudy (turingtest):

AND ^ SYMBOL FOR EXPONENTS!!!!!!!11

OpenStudy (anonymous):

\[9t \sqrt{12^3}+5t \sqrt{128t}-3t \sqrt{48t} \]

OpenStudy (turingtest):

that is way too different from what you have above, I don't know what you mean. If you are SURE the last expression you gave is right, what do you want to do, simplify it?